taylorroop:

Solve 4x − 5 = 6 for x using the change of base formula log base b of y equals log y over log b

taylorroop:

the x-5 is raised

taylorroop:

−4.226 −3.708 5.774 6.292

darkknight:

4^(x-5)=6 log base 4 of 4^(x-5)= log base 4 of 6 x-5=log base 4 of 6 x = log base 4 of 6 plus 5